Suppose a shark propels itself through the water with a forward thrust of 7100 N, resisted by an average drag force of 4200 N. (a) What is the net horizontal force on the shark? (b) What is the net work performed on the shark if it undergoes a horizontal displacement of 30 m? Recall Wnet= Fnetdcosθ. (c) If the shark has an initial velocity of vi = 12m/s, use the work-energy theorem to determine the shark's final velocity after the 30 m displacement described in part (b).
Answer #1

a) net force = 7100 - 4200

= 2900 N

b) work done = F*S

= 2900*30 = 87000 J

c) work done = change in kinetic energy

W = 0.5m(v^2-u^2)

so m is required for calculating the velocity